Using the given bond dissociation energies, calculate ΔH for the reaction, C2H6 (g) + Cl2 (g)  C2H5Cl (g) + HCl (g)

C–C, 348 kJ/mol C–H, 414 kJ/mol Cl–Cl, 242 kJ/mol C–Cl, 327 kJ/mol H–Cl, 431 kJ/mol
Answer #1

Calculation:

  1. Bonds Broken (Reactants):

    • 1 C−H bond in C₂H₆: 414 kJ/mol

    • 1 Cl−Cl bond: 242 kJ/mol
      Total energy absorbed: 414 + 242 = 656 kJ/mol

  2. Bonds Formed (Products):

    • 1 C−Cl bond in C₂H₅Cl: 327 kJ/mol

    • 1 H−Cl bond: 431 kJ/mol
      Total energy released: 327 + 431 = 758 kJ/mol

  3. ΔH Calculation:
    ΔH = (Energy absorbed) − (Energy released)
    ΔH = 656 − 758 = -108 kJ/mol


Answer:
ΔH for the reaction is -108 kJ/mol (exothermic).

answered by: anonymous